# Benchmarking the Spatial Robustness of DNNs via Natural and Adversarial Localized Corruptions

### v1

@@ Abstract
The robustness of DNNs is a crucial factor in safety-critical applications, particularly in complex and dynamic environments where localized corruptions can arise. While previous studies have evaluated the robustness of semantic segmentation (SS) models under whole-image natural or adversarial corruptions, a comprehensive investigation into the spatial robustness of dense vision models under localized corruptions remained underexplored. This paper fills this gap by introducing specialized metrics for benchmarking the spatial robustness of segmentation models, alongside with an evaluation framework to assess the impact of localized corruptions. Furthermore, we uncover the inherent complexity of characterizing worst-case robustness using a single localized adversarial perturbation. To address this, we propose region-aware multi-attack adversarial analysis, a method that enables a deeper understanding of model robustness against adversarial perturbations applied to specific regions. The proposed metrics and analysis were evaluated on 15 segmentation models in driving scenarios, uncovering key insights into the effects of localized corruption in both natural and adversarial forms. The results reveal that models respond to these two types of threats differently; for instance, transformer-based segmentation models demonstrate notable robustness to localized natural corruptions but are highly vulnerable to adversarial ones and vice-versa for CNN-based models. Consequently, we also address the challenge of balancing robustness to both natural and adversarial localized corruptions by means of ensemble models, thereby achieving a broader threat coverage and improved reliability for dense vision tasks.
@@ 1 Introduction
In recent years, deep neural networks (DNNs) have demonstrated remarkable performance across various vision applications, increasing their potential applicability in safety-critical domains such as autonomous systems[19, 27]. Among these scenarios, dense prediction tasks like semantic segmentation (SS) have garnered significant attention. These tasks require models to understand the semantic meaning of each pixel in a given scene, often necessitating the extraction of deep contextual information. In this context, it is crucial to assess the robustness and trustworthiness of SS models, as complex and dynamic environments can give rise to unexpected input variations that challenge the entire system reliability.
@@
Despite this importance, the complexity of DNNs and real-world application scenarios, as in autonomous and assisted driving, has made ensuring and evaluating robustness one of the most challenging problems in modern AI [6, 3, 32, 16]. In the literature, robustness evaluation typically involves analyzing model performance under corrupted input conditions, with two main types of corruption commonly studied: (i) Natural corruptions, which simulate faults or real-world environmental conditions such as sensor noise or adverse weather (e.g., snow, rain) [14]; (ii) Adversarial corruptions, which involve intentionally crafted perturbations designed to manipulate model predictions. While adversarial attacks are often studied as a security threat, they are also a valuable tool for evaluating a model’s robustness under a close approximation of worst-case scenarios, contributing to a sense of verifiability [4, 16].
@@
In recent years, numerous studies have evaluated the robustness of vision models against both natural and adversarial perturbations [14, 32, 17]. However, a comprehensive characterization of their behavior under localized corruptions [18, 29, 30] remains insufficiently explored, particularly for dense prediction tasks such as semantic segmentation. In fact, although poorly explored, localized perturbations are highly relevant to real-world scenarios (e.g., sunlight affecting specific regions of an input image) and need to gather more attention in robustness studies. This gap raises critical questions about the spatial robustness of segmentation models: To what extent do corrupted regions influence predictions in both the corrupted and uncorrupted areas of an image? How effectively can models leverage uncorrupted regions to mitigate the impact of corrupted ones? Addressing these questions is crucial for advancing the robustness and reliability of scene-understanding models in safety-critical environments.
@@
This paper addresses the above challenges by studying the effect of localized corruptions across different models through the lens of both natural transformations and adversarial perturbations. In particular, the work begins by formalizing the concept of localized corruptions and proposing ad-hoc metrics specifically designed to evaluate spatial robustness in different regions of the input for SS models. We then introduce an analysis framework for benchmarking the spatial robustness of SS models against localized natural corruptions, enhancing the understanding of the reliability of DNNs in practical scenarios, where localized corruptions can occur, as in driving environments.
@@
Next, we investigate critical aspects of adversarial attacks within the context of SS tasks, highlighting that identifying a single localized perturbation capable of comprehensively evaluating spatial robustness (e.g., misclassifying as many pixel predictions as possible) is particularly challenging due to the multi-objective nature of the adversarial optimization problem. To address this challenge, we propose region-aware multi-attack adversarial analysis, a method that iteratively generates localized perturbations, each targeting regions previously unaffected by prior attacks. The purpose of this method is to demonstrate that regions initially considered robust (as they are not misclassified by a single localized attack) can indeed be compromised when specific attack configurations are applied. This insight enhances the understanding of worst-case localized perturbations and their impact on model robustness, highlighting the need to understand the extent to which pixels can be perturbed when a perturbation is applied to a specific area.
@@
The proposed framework, considering both localized natural and adversarial corruptions, is then applied in an experimental evaluation to benchmark the spatial robustness of $15$ SS models in driving scenarios [8]. Most interestingly, we observed that lightweight convolution-based SS models are significantly vulnerable to localized natural corruptions, while Transformer-based models exhibit higher robustness. Conversely, the former are more robust to adversarial perturbations, while the latter and highly vulnerable to the same. These findings raise significant concerns about the reliability of SS models when evaluated under worst-case safety-related scenarios.
To address this issue and highlight the importance of achieving a trade-off in terms of spatial robustness between natural transformations and adversarial perturbations, we finally explore the design and configuration of ensemble strategies.
@@
Overall, this study marks a significant step toward understanding localized corruptions and paves the way for developing more robust SS models in practice.
Contribution. In summary, this paper makes the following contribution:
@@
  • We formalize new metrics to evaluate the accuracy of SS models while accounting for their spatial robustness under localized corruptions. Furthermore, we extend existing evaluation approaches to study localized corruption by proposing a flexible framework that allows for quick evaluation of the model spatial robustness.
  • We uncover important insights into the study of localized adversarial perturbations for spatial evaluation in SS models and introduce a new method for generating adversarial perturbations named region-aware multi-attack adversarial analysis.
@@
  • We applied the proposed metrics, framework, and method to analyze the spatial robustness of 15 SS models under natural and adversarial localized perturbations, revealing contrasting behaviors between the two types of corruptions. Finally, we study the trade-off between robustness to natural and adversarial corruptions by means of ensemble models.
@@ 2 Related Work
Deep learning models are widely used across various applications, including safety-critical domains such as autonomous driving. Given their increasing deployment in such scenarios, assessing their robustness is essential to evaluate their reliability and suitability for real-world use. To address these challenges, several benchmarks and robustness analyses have been proposed over the years to assess model reliability against both natural and adversarial corruptions. Notable examples include [14, 10, 17, 31].
Regarding natural corruptions, the objective is to evaluate models’ robustness against perturbations that simulate adverse or potentially out-of-distribution conditions that can be encountered in real-world environments, such as noise, rain, or snow. One of the earliest and most influential studies in this area was conducted by Hendrycks and Dietterich, who introduced a benchmark to assess the resilience of image classification models. This benchmark comprises various natural corruptions, including weather-related disturbances and image formatting artifacts [14]. Since then, these corruptions have been widely adopted and extended in subsequent studies to systematically evaluate and compare the robustness of different models [13, 2, 25]. For instance, Bhojanapalli et al. [2] investigated the robustness of ViTs comparing it to the robustness of a ResNet-50 model under image corruptions, concluding that ViTs models are at least as robust as the CNN take into consideration on the majority of the corruptions. While the majority of studies have focused on classification tasks, fewer works have explored the robustness of models in dense prediction tasks, such as semantic segmentation [31, 17], which have highlighted the need for more comprehensive evaluations in this new domain
Beyond natural corruptions, other works have examined adversarial corruptions [32, 3], which represent a worst-case scenario from a security perspective, where perturbations are deliberately crafted to deceive models. Similar to natural corruptions, several studies have compared the robustness of different architectures against adversarial perturbations [2, 25, 22]. Specifically, research on adversarial robustness in semantic segmentation [37, 1, 17] has underscored the importance of establishing more rigorous benchmarks and developing more effective defense strategies to enhance model resilience in real-world applications.
@@
While most of the aforementioned studies primarily focus on corruptions affecting the entire input, it is crucial to consider the problem of spatial robustness [30, 29, 13]. This aspect is particularly relevant for dense prediction tasks, where model predictions may be influenced by regions beyond the immediate locality of the target area. For instance, in semantic segmentation, pixels that are spatially distant from the region of interest can still impact the model’s decision in unexpected ways and vice versa.
Despite its importance, research on spatial robustness under natural corruptions remains limited. For example, [13] analyzes the robustness of ViTs compared to CNNs under natural and adversarial corruptions concluding that ViTs generally exhibit greater resilience to natural corruptions than CNNs. However, does not specifically address dense prediction tasks. Conversely, most existing studies on spatial robustness have focused on adversarial settings, particularly adversarial patches [26, 5, 20, 23], while only a few works have explored the effects of localized adversarial perturbations [24]. Additionally, there is still a lack of comparative analysis on transformer-based models in both cases.
In this study, we primarily focus on localized adversarial perturbations as they are more directly related to worst-case safety evaluations. While adversarial patches pose a significant security concern, their effects and mitigation strategies have been explored in the literature [28, 13]. Our work aims to bridge this gap by also investigating the unexplored case of natural corruptions, which is particularly relevant for safety assessments. We provide a more comprehensive analysis of spatial robustness in the context of adversarial perturbations, with a specific focus on comparing transformer-based and convolution-based architectures.
@@ 3 Spatial Robustness Analysis
This section first provides some preliminary and background information to theoretically define localized perturbation within the context of SS models. Then, new metrics for spatial robustness are proposed to generalize the classic pixel-wise accuracy score used in semantic segmentation.
@@ 3.1 Background and Formalisms
We consider input images of size $H\times W$ and denote by $\mathcal{I}$ the set of their pixels, which are then individually referred to using the index $i$. A semantic segmentation model designed to classify each pixel into one of $N$ classes is represented by a function $f:x\mapsto\mathbb{R}^{(H\cdot W)\times N}$, which outputs a per-class distribution of scores for each pixel $i$. The vector of output scores for each pixel $i$ is denoted by $f_{i}$ while its component related to the $j$-th class is denoted by $f^{j}_{i}$. The semantic label predicted for pixel $i$ is determined by selecting the class with the highest score for that pixel, i.e., $\operatorname*{argmax}_{j\in\{1,\dots,N\}}f_{i}^{j}(x).$ Finally, the corresponding ground-truth segmentation map for the image $x$ is denoted by $y\in\mathbb{I}^{H\times W}$, where each pixel is assigned its true class label, i.e., $y_{i}\in\{1,\dots,N\}$.
@@
To test the spatial robustness of SS models, we consider corruptions applied to only part of the image. Specifically, following the notation adopted in previous work [24, 29], we define a corrupted area $M$ as a binary mask in the input space $M\in\{0,1\}^{H\times W}$, which indicates where the corruption is applied. For any 2D pixel with index $i$, $M_{i}=1$ denotes the pixel is corrupted, while $M_{i}=0$ denotes that it is not. The corrupted image $x_{M}^{c}$, given a mask $M$ and a corruption function $c(\cdot)$, is computed as:
@@
where $c(\cdot)$ can represent either a natural transformation or an adversarial attack. Additionally, we introduce the ratio of the perturbed area $r=\sum_{i}M_{i}/|M|$ and the non-corrupted region mask, denoted by $\overline{M}$, which is the complement of $M$, i.e., $\overline{M}=1-M$.
@@ 3.2 Metrics for Spatial Robustness
Next, we define metrics for studying the effect of localized corruptions on predictions in different areas of the image, by generalizing both the classic accuracy metrics used for SS and those applied for testing the robustness of image-classification DNNs against corruptions.
@@
Deriving from the analysis of [14] in image classification, given a model $f$ and a set of possible corruptions $\mathcal{C}$ (e.g., Gaussian noise, adversarial perturbation or None to denote no corruptions), the robustness accuracy of an SS model on a dataset $\mathcal{D}$ is defined as:
@@
where $|\mathcal{C}|$ is the number of corruption types, $|\mathcal{I}|$ is the number of pixels in the input, $|\mathcal{D}|$ is the size of the dataset, and $\mathds{1}(\cdot)$ is an operator that returns ’1’ if a condition is ’True’, otherwise ’0’. Following this metric, we can also define the Corruption Error as:
@@
where $A_{\textit{base}}$ refers to the accuracy of a reference model (either $f$ or another one, e.g., AlexNet, as used in [14]). This enables a better understanding of the impact of a given corruption with respect to a reference value.
@@
where $\mathcal{P}_{i}(M)$ is a spatial importance function designed to weigh the significance of misclassifying a pixel $i$ based on the corruption mask $M$. This function plays a critical role in assessing the impact of regions of the input image, allowing the analysis to either focus on regions of higher relevance or those affected by corruption. While, in general, spatial importance functions can be arbitrary defined, in this study we focus on two specific definitions:
  • Non-Corrupted Region Analysis ($A_{\overline{M}}$): this approach focuses on evaluating the spatial robustness of the model by analyzing regions unaffected by corruptions, uniformly weighting all non-corrupted pixels, i.e. $\mathcal{P}_{i}(M)=\overline{M}_{i}/\left|\overline{M}\right|$.
  • Corrupted Region Analysis ($A_{M}$): this approach is the dual of the one above, focusing on corrupted pixels only, i.e., $\mathcal{P}_{i}(M)=M_{i}/|M|$.
The first definition ($A_{\overline{M}}$) is particularly useful for assessing how well the model remains accurate in non-corrupted regions, offering insights into the capability of corruptions to extend beyond the actual corrupted region. Conversely, the second definition ($A_{M}$) focuses on the model’s ability to recover accurate predictions in corrupted regions, potentially leveraging information from non-corrupted regions ($\overline{M}$). This analysis is more meaningful when studying natural corruptions, as adversarial attacks can easily lead to complete misclassification of corrupted regions [29, 24], making robustness assessments in such areas less informative. In our experimental analysis, we highlight that these metrics are expressive for both natural and adversarial corruptions, as they provide measures of resiliency to corruptions within and far from the targeted regions.
It is worth noting that the metrics presented above can be computed for multiple corruption regions: different settings are considered in our experimental evaluation (Section 6). For localized adversarial perturbations, we follow traditional approaches that focus on statically targeting specific areas of the image, such as the center or a corner (e.g.,  [29, 24]). The rationale behind this choice is that these positions allow evaluating the worst- or best-case scenarios of a model’s receptive field [21].
@@
where $M_{r}(x)$ is a mask that determines the corrupted regions, with patches randomly selected for each input $x$ while respecting a ratio $r=\sum_{i}M_{i}/|M|$ of the corrupted area.

### v3

@@ Abstract
The robustness of deep neural networks is a crucial factor in safety-critical applications, particularly in complex and dynamic environments (e.g., medical or driving scenarios) where localized corruptions can arise. While previous studies have evaluated the robustness of semantic segmentation (SS) models under whole-image natural or adversarial corruptions, a comprehensive investigation into the spatial robustness of dense vision models under localized corruptions remains underexplored. This paper fills this gap by introducing novel, region-aware metrics for benchmarking the spatial robustness of segmentation models, along with an evaluation framework to assess the impact of natural localized corruptions. Furthermore, it uncovers the inherent complexity of evaluating worst-case spatial robustness using only a single localized adversarial attack. To address this, the work proposes a region-aware multi-attack adversarial analysis to systematically assess model robustness across specific image regions. The proposed metrics and analysis were exploited to evaluate 14 segmentation models in driving scenarios, uncovering key insights into the effects of localized corruption in both natural and adversarial forms. The results reveal that models respond to these two types of threats differently; for instance, transformer-based segmentation models demonstrate notable robustness to localized natural corruptions but are highly vulnerable to adversarial ones, and vice versa for CNN-based models. Consequently, we also address the challenge of balancing robustness to both natural and adversarial localized corruptions by means of ensemble models, thereby achieving a broader threat coverage and improved reliability for dense vision tasks.
@@ 1 Introduction
In recent years, deep neural networks (DNNs) have demonstrated remarkable performance across various vision applications, increasing their potential applicability in safety-critical domains such as autonomous systems[survey_liu2024survey, PR_QIAN2022108796]. Among these scenarios, dense prediction tasks like semantic segmentation (SS) have garnered significant attention. These tasks require models to understand the semantic meaning of each pixel in a given scene, often necessitating the extraction of deep contextual information. In this context, it is crucial to assess the robustness and trustworthiness of SS models, as complex and dynamic environments can give rise to unexpected input variations that challenge the entire system reliability.
@@
The complexity of DNNs and real-world application scenarios, as in autonomous and assisted driving, has made ensuring and evaluating robustness one of the most challenging problems in modern AI [Carlini017, biggio2018wild, Szegedy14, survey_trust]. In the literature, robustness evaluation typically involves analyzing model performance under corrupted input conditions, with two main types of corruption commonly studied: (i) Natural corruptions, which simulate faults or real-world environmental conditions such as sensor noise or adverse weather (e.g., snow, rain) [hendrycks2018benchmarking]; (ii) Adversarial corruptions (adversarial attacks), which involve intentionally crafted perturbations designed to manipulate model predictions. While adversarial attacks are often studied as a security threat, they are also a valuable tool for evaluating a model’s robustness under a close approximation of worst-case scenarios, contributing to a sense of verifiability [brau2022minimal, survey_trust].
@@
In recent years, numerous studies have evaluated the robustness of vision models against both natural and adversarial perturbations [hendrycks2018benchmarking, Szegedy14, kamann2020benchmarking]. However, a comprehensive characterization of their behavior under localized corruptions [patch_object, rossolini_tnnls_2023, saha_role_2020], which affect only parts of the image, remains insufficiently explored, particularly for dense prediction tasks such as semantic segmentation. Indeed, although poorly explored, localized perturbations are highly relevant to real-world scenarios (e.g., sunlight affecting specific regions of an input image) and need to gather more attention in robustness studies. Exploring localized corruptions allows addressing critical questions about the spatial robustness of semantic segmentation models: To what extent do corrupted regions influence predictions in both the corrupted and uncorrupted areas of an image? How effectively can models leverage uncorrupted regions to mitigate the impact of corrupted ones? Addressing these questions is crucial for advancing the robustness and reliability of scene-understanding models in safety-critical environments.
@@
This paper addresses the above challenges by studying the effect of localized corruptions across different models through the lens of both natural transformations and adversarial perturbations. In particular, the work begins by formalizing the concept of localized corruptions and proposing ad-hoc metrics specifically designed to evaluate spatial robustness in different regions of the input for SS models.
An analysis framework for benchmarking the spatial robustness of SS models against localized natural corruptions is then introduced, enhancing the understanding of the reliability of DNNs in practical scenarios, such as in driving environments, where localized corruptions can occur.
@@
Next, critical aspects of adversarial attacks within the context of SS tasks are investigated, highlighting that identifying a single localized perturbation capable of comprehensively evaluating spatial robustness (e.g., misclassifying as many pixel predictions as possible) is particularly challenging. This is due to the multi-objective nature of the adversarial optimization problem. To address this challenge, we propose the region-aware multi-attack adversarial analysis, a novel evaluation algorithm that iteratively generates localized perturbations, each targeting regions previously unaffected by prior attacks. The idea behind the proposed algorithm builds on the observation that regions initially considered robust (because they are not misclassified by a single localized attack) can, instead, be compromised when repeated runs of the same attack are applied with more targeted configurations. This enables a clearer understanding and evaluation of worst-case localized perturbations, revealing vulnerabilities that may not be sufficiently captured by standard localized attacks.
@@
The proposed framework, considering both localized natural and adversarial corruptions, is then applied in an experimental evaluation to benchmark the spatial robustness of $14$ SS models in driving scenarios [cordts2016cityscapes]. Most interestingly, we observed that addressed convolution-based models are significantly vulnerable to localized natural corruptions, while addressed transformer-based models exhibit higher robustness. Conversely, the opposite trend arises when considering localized adversarial corruptions: convolution-based models appear more robust, whereas transformer-based models appear highly vulnerable due to the heavy exploitation of global attention mechanisms. These findings raise contrastive concerns about the reliability of SS models when evaluated under worst-case safety-related scenarios. To address this, we finally investigate a promising ensembling analysis aimed at achieving a better trade-off in spatial robustness between natural and adversarial perturbations.
@@
Overall, this study marks a significant step toward understanding localized corruptions and paves the way for developing more robust SS models in practice. [^1] To support reproducibility, we provide the code and benchmark implementation used in our experiments at the following repository: https://github.com/GiuliaMP/SpatialRobustnessBench.
In summary, this paper makes the following contributions:
@@
  • It formalizes new metrics to evaluate the accuracy of SS models while accounting for their spatial robustness under localized corruptions. Furthermore, it extends existing approaches to study localized corruption by proposing a flexible framework for evaluating the model spatial robustness.
  • It uncovers important insights and challenges in the study of localized adversarial attacks in SS models and introduces a new evaluation method that provides a more thorough assessment of robustness by iteratively testing multiple attacks.
@@
  • The proposed metrics, frameworks, and methods are exploited to analyze the spatial robustness of 14 SS models on the Cityscape dataset, under natural and adversarial localized perturbations, revealing contrasting behaviors between the two types of corruptions. Finally, the trade-off between natural and adversarial robustness is explored by means of ensemble analysis.
@@ 2 Related Work
Deep learning models are widely used across various applications, including safety-critical domains such as autonomous driving. Given their increasing deployment in such scenarios, assessing their robustness is essential to evaluate their reliability and suitability for real-world use. To address these challenges, several benchmarks and robustness analyses have been proposed over the years to assess model reliability against both natural and adversarial corruptions. Notable examples include [hendrycks2018benchmarking, croce2022interplay, kamann2020benchmarking, schiappa2024robustness].
Regarding natural corruptions, the objective is to evaluate models’ robustness against perturbations that simulate adverse or potentially out-of-distribution conditions that can be encountered in real-world environments, such as noise, rain, or snow. One of the earliest and most influential studies in this area was conducted by Hendrycks and Dietterich, who introduced a benchmark to assess the resilience of image classification models. This benchmark comprises various natural corruptions, including weather-related disturbances and image formatting artifacts [hendrycks2018benchmarking]. Since then, these corruptions have been widely adopted and extended in subsequent studies to systematically evaluate and compare the robustness of different models [gu2022vision, bhojanapalli2021understanding, naseer2021intriguing]. For instance, Bhojanapalli et al. [bhojanapalli2021understanding] investigated the robustness of ViTs comparing it to the robustness of a ResNet-50 model under image corruptions, concluding that ViTs models are at least as robust as the CNN taking into consideration on the majority of the corruptions. While the majority of studies have focused on classification tasks, fewer works have explored the robustness of models in dense prediction tasks, such as semantic segmentation [schiappa2024robustness, kamann2020benchmarking], which have highlighted the need for more comprehensive evaluations in this new domain.
Beyond natural corruptions, other works have examined adversarial corruptions [Szegedy14, biggio2018wild], which represent a worst-case scenario from a security perspective, where perturbations are deliberately crafted to deceive models. Similar to natural corruptions, several studies have compared the robustness of different architectures against adversarial perturbations [bhojanapalli2021understanding, naseer2021intriguing, mahmood2021robustness]. Specifically, research on adversarial robustness in semantic segmentation [yuan2024towards, arnab_robustness_nodate, kamann2020benchmarking] has underscored the importance of establishing more rigorous benchmarks and developing more effective defense strategies to enhance models’ robustness in real-world applications.
@@
While most of the aforementioned studies primarily focus on corruptions affecting the entire input, it is crucial to consider the problem of spatial robustness [saha_role_2020, rossolini_tnnls_2023, gu2022vision]. This aspect is particularly relevant for dense prediction tasks, where model predictions may be influenced by regions beyond the immediate locality of the target area. For instance, since semantic segmentation performance is usually evaluated at a pixel level, defining a robustness analysis for such a task opens the possibility to study the impact of perturbed regions on clean regions, and vice versa.
Despite its importance, research on spatial robustness under natural corruptions remains limited. For example, [gu2022vision] analyzes the robustness of ViTs compared to CNNs under natural and adversarial corruptions, concluding that ViTs generally exhibit greater resilience to natural corruptions than CNNs. However, does not specifically address dense prediction tasks. Conversely, most existing studies on spatial robustness have focused on physically-realizable adversarial settings, e.g., adversarial patches [pintor_imagenetpatch, brown_adversarial_2018, liu2025_patch, metzen2021meta, rossolini_iccps24], while only a few works have explored the effects of single localized adversarial perturbations [nakka_indirect_2020], which instead reveal intrinsic and human-imperceptible weaknesses of the model. Additionally, there is still a lack of comparative analysis on transformer-based models in both cases.
Differently from these works, this paper shows that, from an adversarial perspective, a single attack is not sufficient to properly represent the true spatial robustness of dense prediction models. To this end, it proposes an ad hoc evaluation analysis that iteratively adjusts the fooling regions across multiple attacks in the same area, enabling the assessment of regions that conventional attacks might leave unaffected. Furthermore, to address the lack of analysis on natural localized robustness, an aspect particularly relevant for safety assessments, this work introduces a schematic framework and ad hoc metrics to benchmark models in these settings.
To the best of our knowledge, this work provides the first comprehensive analysis of spatial robustness in the context of both natural and adversarial perturbations for semantic segmentation. This study allowed uncovering important experimental insights when comparing transformer-based and convolution-based architectures and proposing a promising ensembling strategy.
@@ 3 Spatial Robustness Analysis
This section first provides some preliminary and background information to theoretically define localized corruptions within the context of SS models. Then, new metrics for spatial robustness are proposed to generalize the classic pixel-wise accuracy score used in semantic segmentation.
@@ 3.1 Background and Formalisms
We consider input images of size $H\times W$ and denote by $\mathcal{I}$ the set of the image’s pixels, which are then individually referred to using the index $i$. A semantic segmentation model designed to classify each pixel into one of $N$ classes is represented by a function $f:x\mapsto\mathbb{R}^{(H\cdot W)\times N}$, which outputs a per-class distribution of scores for each pixel $i$. The vector of output scores for each pixel $i$ is denoted by $f_{i}$ while its component related to the $j$-th class is denoted by $f^{j}_{i}$. The semantic label predicted for pixel $i$ is determined by selecting the class with the highest score for that pixel, i.e., $\operatorname*{argmax}_{j\in\{1,\dots,N\}}f_{i}^{j}(x).$ Finally, the corresponding ground-truth segmentation map for the image $x$ is denoted by $y\in\mathbb{N}^{H\times W}$, where each pixel is assigned its true class label, i.e., $y_{i}\in\{1,\dots,N\}$.
@@
To test the spatial robustness of SS models, we consider corruptions applied to only part of the image. Specifically, following the notation adopted in previous work [nakka_indirect_2020, rossolini_tnnls_2023], we define a corrupted area $M$ as a binary mask in the input space $M\in\{0,1\}^{H\times W}$, which indicates where the corruption is applied. For any 2D pixel with index $i$, $M_{i}=1$ denotes the pixel is corrupted, while $M_{i}=0$ denotes that it is not. The corrupted image $x_{M}^{c}$, given a mask $M$ and a corruption function $c(\cdot)$, is computed as:
@@
where $c(\cdot)$ can represent either a natural transformation or an adversarial attack. Additionally, we introduce the ratio of the perturbed area as $r=S/(H\cdot W)$, where $\mathcal{S}$ is the number of corrupted pixels, computed from $M$ as $\mathcal{S}=\sum_{i=1}^{H\cdot W}M_{i}$. It follows that the number $\overline{\mathcal{S}}$ of non-corrupted pixels is $\overline{\mathcal{S}}=\sum_{i=1}^{H\cdot W}1-M_{i}$.
@@ 3.2 Metrics for Spatial Robustness
We define metrics for studying the effect of localized corruptions on predictions in different areas of the image, by generalizing both the classic accuracy metrics used for SS and those applied for testing the robustness of image-classification DNNs against corruptions.
@@
Deriving from the analysis of [hendrycks2018benchmarking] in image classification, given a model $f$ and a set of possible corruptions $\mathcal{C}$ (e.g., Gaussian noise, adversarial perturbation or $\{\emptyset\}$ to denote no corruptions), the robustness accuracy of an SS model on a dataset $\mathcal{D}$ is defined as:
@@
where $|\mathcal{C}|$ is the number of corruption types, $|\mathcal{I}|$ is the number of pixels in the input, $|\mathcal{D}|$ is the size of the dataset, and $\mathds{1}(\cdot)$ is an operator that returns ‘1’ if a condition is ‘True’, otherwise ‘0’. Following this metric, we can also define the Corruption Error as:
@@
where $A_{\textit{base}}$ refers to the accuracy of a reference model (either $f$ or another one, e.g., AlexNet, as used in [hendrycks2018benchmarking]). This enables a better understanding of the impact of a given corruption with respect to a reference value.
@@
where $\mathcal{P}_{i}(M)$ is a spatial importance function designed to weigh the significance of misclassifying a pixel $i$ based on the corruption mask $M$. This function plays a critical role in assessing the impact of regions of the input image, allowing the analysis to either focus on regions of higher relevance or those affected by corruption. While, in general, spatial importance functions can be arbitrarily defined, in this study we focus on two specific definitions:
  • Non-Corrupted Region Analysis ($A_{\overline{M}}$): this approach focuses on evaluating the spatial robustness of the model by analyzing regions unaffected by corruptions, uniformly weighting all non-corrupted pixels, i.e. $\mathcal{P}_{i}(M)=(1-M_{i})/\overline{\mathcal{S}}$.
  • Corrupted Region Analysis ($A_{M}$): this approach is the dual of the one above, focusing on corrupted pixels only, i.e., $\mathcal{P}_{i}(M)=M_{i}/\mathcal{S}$.
Notice that both $A_{\overline{M}}$ and $A_{M}$ are normalized by the number of pixels in their respective regions, ensuring fairness in evaluating and comparing each region. Such localized metrics reveal whether the corruption affects predictions beyond its regions, or whether the model can leverage clean informations from outside the perturbed areas to correctly classify the perturbed pixels.
The first definition ($A_{\overline{M}}$) is particularly useful for assessing how well the model remains accurate in non-corrupted regions, offering insights into the capability of corruptions to extend beyond the actual corrupted region. Conversely, the second definition ($A_{M}$) focuses on the model’s ability to recover accurate predictions in corrupted regions, potentially leveraging information from non-corrupted regions. This analysis is more meaningful when studying natural corruptions, as adversarial attacks can easily lead to complete misclassification of corrupted regions [rossolini_tnnls_2023, nakka_indirect_2020], making robustness assessments in such areas less informative. It is worth noting that the metrics presented above can be computed for multiple corruption regions: different settings are considered in our experimental evaluation (Section 6). For localized adversarial perturbations, we follow traditional approaches that focus on statically targeting specific areas of the image, such as the center or a corner (e.g.,  [rossolini_tnnls_2023, nakka_indirect_2020]). The rationale behind this choice is that these positions allow evaluating the worst- or best-case scenarios of a model’s receptive field [luo2016understanding].
@@
where $M_{r}(x)$ is a mask that determines the corrupted regions, with patches randomly selected for each input $x$ while respecting the ratio $r=S/(H\cdot W)$ of the corrupted area.